# spacing=nonfrench

### 04 Mar 2026 v1

@@ 3.4. The main theorem
We have thus constructed $\Gamma_{f}$, a height-$1$ connected $p$-divisible group over $\mathcal{O}_{L}$, whose followed path is summarized below:
$\operatorname{\mathbf{Rep}}_{\mathbb{Z}_{p}}^{\mathrm{cris},1}(G_{L})$$\operatorname{\mathbf{BT}}^{\mathrm{f}}_{\mathfrak{S}}(\varphi)$$S\textbf{-}\mathbf{Win}/\mathcal{O}_{L}$$(p\textbf{-}\mathbf{Div}/\mathcal{O}_{L})^{\mathrm{c}}$$T_{f}$$\mathfrak{M}_{f}$$W_{f}$$\Gamma_{f}$}
and ensure that $f$ is an endomorphism of each of these objects. It follows that Lubin’s conjecture holds in our setting.
@@
The group law of $\Gamma_{f}$ induces a comultiplication on its Hopf algebra; this is a map $\mathcal{O}_{L}[\![T]\!]\rightarrow\mathcal{O}_{L}[\![X,Y]\!]$ given by sending $T$ to $F(X,Y)$, where $F(X,Y)$ is a height-$1$ formal group defined over $\mathcal{O}_{L}$. By construction, $f$ is an endomorphism of $F$, and so $F$ is defined over $\mathcal{O}_{L}\cap K=\mathcal{O}_{K}$ by the uniqueness property stated in §1.1. As $u$ is defined over $\mathcal{O}_{K}$ and commutes with $f$, it is necessarily an $\mathcal{O}_{K}$-endomorphism of $F$ as well. ∎

### 22 Apr 2026 v2

@@ 3.4. The main theorem
We have thus constructed $\Gamma_{f}$, a height-one connected $p$-divisible group defined over $\mathcal{O}_{L}$, whose followed path is summarized below:
$\operatorname{\mathbf{Rep}}_{\mathbb{Z}_{p}}^{\mathrm{cris},\mathrm{fgl}}(G_{L})$$\operatorname{\mathbf{BT}}^{\mathrm{f}}_{\mathfrak{S}}(\varphi)$$T_{f}$$\mathfrak{M}_{f}$$(p\textbf{-}\mathbf{Div}/\mathcal{O}_{L})^{\mathrm{c}}$$S\textbf{-}\mathbf{Win}/\mathcal{O}_{L}$$\Gamma_{F}$$W_{f}$}
and ensure that $f$ is an endomorphism of all of these objects. It follows that Lubin’s conjecture holds in our setting.
@@
The group law of $\Gamma_{f}$ induces a comultiplication on its Hopf algebra; this is a map $\mathcal{O}_{L}[\![T]\!]\rightarrow\mathcal{O}_{L}[\![X,Y]\!]$ given by sending $T$ to $F(X,Y)$ a height-one formal group defined over $\mathcal{O}_{L}$. By construction $f$ is an endomorphism of $\Gamma_{f}$ and therefore is an endomorphism of $F$. By the uniqueness property stated in §1.1, we deduce that $F$ is defined over $\mathcal{O}_{L}\cap K=\mathcal{O}_{K}$. Finally, as $u$ is defined over $\mathcal{O}_{K}$ and commutes with $f$, it is necessarily an $\mathcal{O}_{K}$-endomorphism of $F$ as well. ∎